Multiplying and Dividing Radicals
Simplify.
1) 2+√3÷√3
2) √11+ 6 ÷ √27
3) 2√7x²y÷ -3√18x³y
4) 48√42x² ÷ 8 √6x

Answer by stanbon(57219)

1) (2+√3)÷√3 = [sqrt(3)*(2+sqrt(3)]/3 = (3+2sqrt(3))/3
---------------------------------------------------------------
2) (√11+ 6) ÷ √27 = [sqrt(3)(sqrt(11) + 6)/9= [6sqrt(3)+sqrt(33))/9
----------------------------------------------------------------------
3) 2√7x²y÷ -3√18x³y = [(2sqrt(7)x^2y)/(-9sqrt(2)x^3y)][sqrt(2)/sqrt(2)]
--------------------------------------------------------------------------
= [2sqrt(14)]/[-18x] = [-sqrt(14)]/[9x]
----------------------------------------------
4) 48√42x² ÷ 8 √6x = 6sqrt(7)x
=====================================